OpenStudy (anonymous):

Input the equation of the given line in standard form. The line with m = 2/3 and passing through (1, 1)

OpenStudy (anonymous):

Use from this !;) Y2-Y=m(X2-X) In there : X2=1 Y2=1 Y and X are in standard form !:)

OpenStudy (anonymous):

so its y=x?

OpenStudy (anonymous):

im a bit confused.

OpenStudy (anonymous):

look: 1-y=2/3(1-x) 1-y=2/3-2/3x -y=-1+2/3-2/3x y=1-2/3+2/3x y=1/3+2/3x Got it ?:)

OpenStudy (anonymous):

i got it Thanks for your help. :)

OpenStudy (anonymous):

You re welcome !